Two springs are hooked together end to end. When a 4.8-kg brick is suspended from one end of the combination, the combination stretches 0.15 m beyond its relaxed length. What is the spring constant of the combination? If the top spring stretches 0.10 m, what is the spring constant of each spring?
Answer #1

when two springs are in series

requvalent spring constant 1/keq=1/k1+1/k2

keq=k1k2/k1+k2

total stretch is 0.15m

f=kx=mg

keq=4.8*9.8/0.15=313.6 N/m

if top spring stretches 0.10 m

k1 = mg / l

k1=4.8*9.8/0.10

k1=470.4 N/m

1/k2=1/313.6-1/470.4

k2= 940.8 N/m